# Geometrical properties of strained and twisted moiré heterostructures

### v1

@@ III.1 Hexagonal homobilayers
From the moiré vectors one can describe the geometrical properties of the moiré pattern by two parameters: their relative angle $\beta$ and their relative length [kogl2023moire, escudero2024designing]. For moiré patterns arising from only a relative rotation of the honeycomb lattices (i.e., with no strain), $\beta=120^{\circ}$ is always the same as the angle between the reciprocal vectors $\mathbf{b}_{i}$. This changes in the presence of strain due to the lattice deformation. The angle $\beta$ can be obtained through the relation $\mathbf{G}_{1}\cdot\mathbf{G}_{2}=\left|\mathbf{G}_{1}\right|\left|\mathbf{G}_{2}\right|\cos\beta$. Using that $\mathbf{G}_{i}=\mathbf{T}\mathbf{b}_{i}$ and the property [gurtin2010mechanics] $\mathbf{T}\mathbf{b}_{1}\cdot\mathbf{T}\mathbf{b}_{2}=\left(\mathbf{T}^{\mathrm{T}}\mathbf{T}\right)\mathbf{b}_{1}\cdot\mathbf{b}_{2}$, where $\mathbf{T}^{\mathrm{T}}$ is the transpose of $\mathbf{T}$, it follows that $\beta$ is uniquely determined by the symmetric transformation $\mathbf{F}=\mathbf{T}^{\mathrm{T}}\mathbf{T}$ acting on the reciprocal vectors [escudero2024designing]
@@
The transformation by $\mathbf{F}_{0}$ is a spherical tensor (i.e., a tensor proportional to the identity matrix $\mathbb{I}$), so rotations alone do not change the angle of the moiré vectors. However, in the presence of strain the non-zero transformation $\mathbf{F}_{\epsilon}$ can modify the angle $\beta$ in Eq. (47), and thus the geometry of the moiré patterns. Note that the expression of $\mathbf{F}_{\epsilon}$ depends the interplay between twist and strain; in particular, $\mathbf{F}_{\epsilon}$ is not a spherical tensor only for strain configurations that deform the lattices unit cell (in other words, for non-spherical strain tensors $\mathcal{E}$). The last term in $\mathbf{F}_{\epsilon}$ describes the possibility of obtaining moiré patterns purely by strain [san2012non].
Significant modifications of the moiré vectors by the strain can be archived even at very small strain magnitudes, which may only barely distort the underlying lattices [kogl2023moire, escudero2024designing]. This is because the observed moiré generally acts as a magnifying glass that enhances any small lattice mismatch between the layers [cosma2014moire]. One can see this in a simple example of both twist and uniaxial heterostrain with magnitude $\epsilon$ along the $x$ direction. In such case the angle between the moiré vectors, to leading order in $\theta$ and $\epsilon_{u}$, is given by [escudero2024designing]
@@
The dependence $\sim\epsilon_{u}/\theta$ reflects that the smaller the twist angle, the smaller the strain magnitude needed to modify the moiré geometry. As a result, for low twist angles and experimentally relevant values $\epsilon\lesssim 10\%$, one can in principle vary the angle $\beta$ to any value between 0 and $180^{\circ}$. By comparison, such strain range only modifies the angle between the lattice vectors of the honeycomb lattices by, at most, a few degrees [naumis2017electronic]. Superlattice configurations thus offer a rich and accessible experimental platform in which one can vary, or effectively design, practically any desired moiré geometry. Next, we describe the possible moiré geometries under the three types of strain introduced in Section II.2.
@@ III.1.1 Uniaxial strain
Here we have used that $\mathbf{F}\mathbf{b}_{1}\cdot\mathbf{b}_{2}=\mathbf{F}\mathbf{b}_{2}\cdot\mathbf{b}_{1}$, since $\mathbf{F}$ is a symmetric tensor [gurtin2010mechanics]. Using Eq. (53) and solving Eq. (55) for the non-zero strain case gives [escudero2024designing]
@@
Note that $\epsilon_{eq}$ is not symmetric under the transformation $\phi\rightarrow\phi+\pi/3$ because, and discussed before, the particular moiré vectors given by Eq. (45) do not take into account all the symmetrical solutions obtained by appropriately changing the moiré vectors construction [escudero2024designing]. The hexagonal symmetry is restored by taking different set of strained reciprocal vectors. For most proposes, it is sufficient to consider $\epsilon_{eq}$ above, and then generalize the results by taking all the missing solutions given by translations $\phi\rightarrow\phi\pm\pi/3$. Keeping that in mind, by solving the angle equation (47) one can find that the strain parameters that yield an angle $\beta$ between the moiré vectors read [escudero2024designing]
@@
Figure 3: Examples of moiré patterns in two equal hexagonal layers with relative twist $\theta=2^{\circ}$ and uniaxial heterostrain. All cases shown correspond to equal length moiré vectors (Section III.1.1). Panel (a) shows three moiré pattern corresponding to (from left to right): $\epsilon=0$ (no strain), $\epsilon\approx 1.64\%$, $\phi\approx-9.4^{\circ}$, $\epsilon\approx 4.44\%$, $\phi\approx-0.9^{\circ}$ and $\epsilon\approx-2.28\%$, $\phi\approx-22.7^{\circ}$. The relative angles between the real space moiré vectors are, respectively, $\beta_{R}=60^{\circ},90^{\circ},140^{\circ},30^{\circ}$. In each case the Wiger-Seitz cell of the superlattice is shown in white. The small inset below each panel show the strain magnitude (in a scale of $5\%$) and the strain direction relative to the $x$ axis. Panel (b) shows the evolution of the Wigner-Seitz cell and the repeated moiré pattern within it, for moiré vector angles (from left to right) $\beta_{R}=40^{\circ},60^{\circ},80^{\circ},90^{\circ},100^{\circ},120^{\circ},140^{\circ}$. For each case, the bar underneath indicates the strain magnitude (thick line) in a scale of $5\%$ (thin line). Adapted under the terms of the $C$ BY license from Ref. [escudero2024designing]. Copyright (2024) by the American Physical Society.
Some moiré patterns obtained from uniaxial strain can be seen in Fig. 3. Due to the magnifying effect of the moiré, small changes in the strain parameters can lead to significant changes in the resulting superlattice geometry. Consequently, it becomes imperative to have a precise control of the strain parameters in order to design particular moiré geometries [pena2023moire] (see Section IV). Conversely, such strong dependence on the strain parameters indirectly explains the plethora of different moiré geometries seen in experiments [alden2013strain, bai2020excitons, hesp2024cryogenic], particularly in those where strain effects unavoidable arise during the superlattice assemble [kazmierczak2021strain].
@@ III.1.2 Shear strain
The last term, which depends on the interplay between twist and strain, is not a spherical tensor so it changes the moiré geometry. In other words, the moiré patterns with both twist and shear strain are not hexagonal [kogl2023moire, escudero2024designing]. The possible strained configurations are similar to those described in the previous section for uniaxial heterostrain (this analogy becomes exact with a combination of shear and biaxial strain, see Section III.1.4).
@@ III.1.3 Biaxial strain
As this is, as expected, a spherical tensor, a biaxial strain does not change the moiré geometry. Its effect is to only modify the orientation and length of the moiré vectors. However, both changes are important and can actually significantly influence the geometric and electronic properties of the system [escudero2024designing].
@@
The moiré periodicity for a given twist angle $\theta$ decreases as the strain increases. The above condition does not, however, guarantee that both moiré patterns (with and without strain) would be aligned because their orientation may differ. This can be important in superlattice configurations made up of three (or more) layers with a lattice mismatch, and where only two layers are relatively rotated. In general, the angle $\alpha$ of the moiré pattern orientation, with respect to the non-strain case, reads [escudero2024designing]
@@
A relevant example of the importance of matching the periodicity and orientation condition occurs in heterostructures of TBG/hBN, in which hBN acts as a substrate of TBG [yankowitz2012emergence, xue2011scanning, decker2011local]. This situation can be described by considering three equal graphene monolayers, but with a biaxial strain of magnitude $\epsilon_{b}\sim 1-a_{g}/a_{h}=0.016$ in the bottom layer [cea2020band], in order to account for the lattice mismatch between graphene ($a_{g}=2.46\;\textrm{Å}$) and hBN ($a_{h}=2.50\;\textrm{Å}$). Since the moiré pattern remains hexagonal with twist and biaxial strain, the orientation match between the graphene/hBN moiré and the TBG moiré occurs when the angle $\alpha$ in Eq. (64) is $60^{\circ}$ (or integers values of it), so that $\cos\alpha=\pm 1/2$. Combining this with the equal periodicity condition of Eq. (63), one finds the moiré alignment conditions [long2022atomistic, long2023electronic, cea2020band, escudero2024designing]
@@ III.1.4 Shear and biaxial strain
The possible geometries of the strained moiré pattern, as given by the transformation (68), are analogous to those that result from uniaxial heterostrain. One can actually state this analogy precisely by noting that both cases are equivalent by the correspondence [escudero2024designing]
@@
This allows one to readily obtain the geometrical properties of the strained moiré patterns due to shear and biaxial strain from those previously analyzed for uniaxial heterostrain (Section III.1.1). In particular, one can directly replace the above correspondence in Eqs. (57) and (58) and obtain the needed combination of shear and biaxial strain in order to obtain equal-length moiré vectors with an angle $\beta$ between them. It should be noted that this analogy between different kinds of strain is exact, it does not just alludes to an equivalence of periodicity or orientation. Indeed, the above correspondence leads to equivalent uniaxial and shear+biaxial strain tensors, so that each lattice is exactly strained in the same way. Thus, one can think of uniaxial heterostrain as a particular combination of biaxial and shear strain [kogl2023moire, escudero2024designing]. As the Poisson’s ratio is generally fixed by the properties of the material, this analogy necessarily fixes the ratio between shear and biaxial strain magnitude
@@
Figure 4: Variation of the real space moiré vectors lengths $A^{\prime}_{1}=\left|\mathbf{G}_{1}^{R}\right|$ and $A^{\prime}_{2}=\left|\mathbf{G}_{2}^{R}\right|$, as a function of the twist angle and the strain magnitude, for the cases of two homobilayer ($\mathrm{WSe_{2}-WSe_{2}}$) and heterobilayer ($\mathrm{WSe_{2}}-\mathrm{MoSe_{2}}$) hexagonal lattices. The configuration is such that the top layers is only twisted and the bottom layer is only strained. Panel (a) shows the moiré lengths $A^{\prime}_{1},A^{\prime}_{2}$ (equal) under biaxial strain magnitude $\epsilon_{b}$; dotted ellipses indicate the regions of moiré lenghts $A^{\prime}_{1},A^{\prime}_{2}=10\,\mathrm{nm}.$ Panel (b) shows the moiré lengths $A^{\prime}_{1}$ (top) and $A^{\prime}_{2}$ (bottom) under uniaxial strain magnitude $\epsilon_{u}$ at fixed direction $\phi=0^{\circ}$. The spots labeled as I, II and III indicate the unidimensional, square and hexagonal moiré geometries, respectively (cf. Section III.4). The points $\nu_{1}$ and $\nu_{2}$ in the heterobilayer case indicate the vertexes of the divergence curves (unidimensional patterns). Adapted under the terms of the $C$ BY license from Ref. [kogl2023moire]. Copyright (2023) The author(s).
@@ III.2 Hexagonal heterobilayers
The analysis of the previous section can be directly generalized to the case of two hexagonal layers with different lattice constant. We refer to these as heterobilayers. Since then the non-deformed lattice vectors in each lattice are different, one cannot determine the moiré vectors through a transformation $\mathbf{T}$ as in Eq. (46). However, one can still define a generalized transformation $\mathbf{T}_{h}$ by accounting the lattice mismatch with a biaxial strain $\mathcal{E}_{b}=\mathbb{I}\epsilon_{b}$ [cosma2014moire, kogl2023moire]. Since the moiré pattern only depends on relative deformations, it is convenient to introduce a biaxial strain tensor $\pm\mathcal{E}_{b}/2$ so that the moiré vectors in heterobilayers can be determined as $\mathbf{G}_{i}=\mathbf{T}_{h}\mathbf{b}_{i}$, where
@@
Figure 4 shows the variation in the length of the real space moiré vectors $A^{\prime}_{1},A^{\prime}_{2}=\left|\mathbf{G}_{1}^{R}\right|,\left|\mathbf{G}_{2}^{R}\right|$, as a function of the twist angle and the strain magnitude (biaxial or uniaxial), for the cases of hexagonal homobilayers and heterobilayers. The results correspond to $\mathrm{WSe_{2}}$ and $\mathrm{MoSe_{2}-WSe_{2}}$ homobilayers and heterobilayers, with lattice constant $a_{\mathrm{MoSe_{2}}}=0.329\,\mathrm{nm}$ and $0.4\%$ smaller for $a_{\mathrm{WSe_{2}}}$, and Poisson’s ratios $\nu_{\mathrm{MoSe_{2}}}=0.23$, $\nu_{\mathrm{WSe_{2}}}=0.19$. The configuration is not symmetric: the upper layer is only twisted while the lower layer is only strained.
@@
Figure 5: Variation of the real-space angle $\beta_{R}$ between the moiré vectors (top) and the moiré area $M=\left|\mathbf{G}_{1}^{R}\times\mathbf{G}_{2}^{R}\right|$ (bottom), as a function of the twist angle and the strain magnitude, for the cases of two homobilayer ($\mathrm{WSe_{2}-WSe_{2}}$) and heterobilayer ($\mathrm{WSe_{2}}-\mathrm{MoSe_{2}}$) hexagonal lattices (same configuration as in Figure 4). Panel (a) shows the case of uniaxial strain with magnitude $\epsilon_{u}$ at fixed direction $\phi=0^{\circ}$. Panel (b) shows the case of shear strain with magnitude $\epsilon_{s}$ at fixed direction $\phi=0^{\circ}$. In both cases, the spots labeled as I, II and III indicate the unidimensional, square and hexagonal moiré geometries (see also Section III.4). Adapted under the terms of the $C$ BY license from Ref. [kogl2023moire]. Copyright (2023) The author(s).

### v2

@@ III.1 Hexagonal homobilayers
From the moiré vectors one can describe the geometrical properties of the moiré pattern by two parameters: their relative angle $\beta$ and their relative length [60, 34]. For moiré patterns arising from only a relative rotation of the honeycomb lattices (i.e., with no strain), $\beta=120^{\circ}$ is always the same as the angle between the reciprocal vectors $\mathbf{b}_{i}$. This changes in the presence of strain due to the lattice deformation. The angle $\beta$ can be obtained through the relation $\mathbf{G}_{1}\cdot\mathbf{G}_{2}=\left|\mathbf{G}_{1}\right|\left|\mathbf{G}_{2}\right|\cos\beta$. Using that $\mathbf{G}_{i}=\mathbf{T}\mathbf{b}_{i}$ and the property [43] $\mathbf{T}\mathbf{b}_{1}\cdot\mathbf{T}\mathbf{b}_{2}=\left(\mathbf{T}^{\mathrm{T}}\mathbf{T}\right)\mathbf{b}_{1}\cdot\mathbf{b}_{2}$, where $\mathbf{T}^{\mathrm{T}}$ is the transpose of $\mathbf{T}$, it follows that $\beta$ is uniquely determined by the symmetric transformation $\mathbf{F}=\mathbf{T}^{\mathrm{T}}\mathbf{T}$ acting on the reciprocal vectors [34]
@@
The transformation by $\mathbf{F}_{0}$ is a spherical tensor (i.e., a tensor proportional to the identity matrix $\mathbb{I}$), so rotations alone do not change the angle of the moiré vectors. However, in the presence of strain the non-zero transformation $\mathbf{F}_{\epsilon}$ can modify the angle $\beta$ in Eq. (47), and thus the geometry of the moiré patterns. Note that the expression of $\mathbf{F}_{\epsilon}$ depends the interplay between twist and strain; in particular, $\mathbf{F}_{\epsilon}$ is not a spherical tensor only for strain configurations that deform the lattices unit cell (in other words, for non-spherical strain tensors $\mathcal{E}$). The last term in $\mathbf{F}_{\epsilon}$ describes the possibility of obtaining moiré patterns purely by strain [98].
Significant modifications of the moiré vectors by the strain can be archived even at very small strain magnitudes, which may only barely distort the underlying lattices [60, 34]. This is because the observed moiré generally acts as a magnifying glass that enhances any small lattice mismatch between the layers [27]. One can see this in a simple example of both twist and uniaxial heterostrain with magnitude $\epsilon$ along the $x$ direction. In such case the angle between the moiré vectors, to leading order in $\theta$ and $\epsilon_{u}$, is given by [34]
@@
The dependence $\sim\epsilon_{u}/\theta$ reflects that the smaller the twist angle, the smaller the strain magnitude needed to modify the moiré geometry. As a result, for low twist angles and experimentally relevant values $\epsilon\lesssim 10\%$, one can in principle vary the angle $\beta$ to any value between 0 and $180^{\circ}$. By comparison, such strain range only modifies the angle between the lattice vectors of the honeycomb lattices by, at most, a few degrees [86]. Superlattice configurations thus offer a rich and accessible experimental platform in which one can vary, or effectively design, practically any desired moiré geometry. Next, we describe the possible moiré geometries under the three types of strain introduced in Section II.2.
@@ III.1.1 Uniaxial strain
Here we have used that $\mathbf{F}\mathbf{b}_{1}\cdot\mathbf{b}_{2}=\mathbf{F}\mathbf{b}_{2}\cdot\mathbf{b}_{1}$, since $\mathbf{F}$ is a symmetric tensor [43]. Using Eq. (53) and solving Eq. (55) for the non-zero strain case gives [34]
@@
Note that $\epsilon_{eq}$ is not symmetric under the transformation $\phi\rightarrow\phi+\pi/3$ because, and discussed before, the particular moiré vectors given by Eq. (45) do not take into account all the symmetrical solutions obtained by appropriately changing the moiré vectors construction [34]. The hexagonal symmetry is restored by taking different set of strained reciprocal vectors. For most proposes, it is sufficient to consider $\epsilon_{eq}$ above, and then generalize the results by taking all the missing solutions given by translations $\phi\rightarrow\phi\pm\pi/3$. Keeping that in mind, by solving the angle equation (47) one can find that the strain parameters that yield an angle $\beta$ between the moiré vectors read [34]
@@
Figure 3: Examples of moiré patterns in two equal hexagonal layers with relative twist $\theta=2^{\circ}$ and uniaxial heterostrain. All cases shown correspond to equal length moiré vectors (Section III.1.1). Panel (a) shows three moiré pattern corresponding to (from left to right): $\epsilon=0$ (no strain), $\epsilon\approx 1.64\%$, $\phi\approx-9.4^{\circ}$, $\epsilon\approx 4.44\%$, $\phi\approx-0.9^{\circ}$ and $\epsilon\approx-2.28\%$, $\phi\approx-22.7^{\circ}$. The relative angles between the real space moiré vectors are, respectively, $\beta_{R}=60^{\circ},90^{\circ},140^{\circ},30^{\circ}$. In each case the Wiger-Seitz cell of the superlattice is shown in white. The small inset below each panel show the strain magnitude (in a scale of $5\%$) and the strain direction relative to the $x$ axis. Panel (b) shows the evolution of the Wigner-Seitz cell and the repeated moiré pattern within it, for moiré vector angles (from left to right) $\beta_{R}=40^{\circ},60^{\circ},80^{\circ},90^{\circ},100^{\circ},120^{\circ},140^{\circ}$. For each case, the bar underneath indicates the strain magnitude (thick line) in a scale of $5\%$ (thin line). Adapted under the terms of the $C$ BY license from Ref. [34]. Copyright (2024) by the American Physical Society.
Some moiré patterns obtained from uniaxial strain can be seen in Fig. 3. Due to the magnifying effect of the moiré, small changes in the strain parameters can lead to significant changes in the resulting superlattice geometry. Consequently, it becomes imperative to have a precise control of the strain parameters in order to design particular moiré geometries [94] (see Section IV). Conversely, such strong dependence on the strain parameters indirectly explains the plethora of different moiré geometries seen in experiments [1, 7, 49], particularly in those where strain effects unavoidable arise during the superlattice assemble [56].
@@ III.1.2 Shear strain
The last term, which depends on the interplay between twist and strain, is not a spherical tensor so it changes the moiré geometry. In other words, the moiré patterns with both twist and shear strain are not hexagonal [60, 34]. The possible strained configurations are similar to those described in the previous section for uniaxial heterostrain (this analogy becomes exact with a combination of shear and biaxial strain, see Section III.1.4).
@@ III.1.3 Biaxial strain
As this is, as expected, a spherical tensor, a biaxial strain does not change the moiré geometry. Its effect is to only modify the orientation and length of the moiré vectors. However, both changes are important and can actually significantly influence the geometric and electronic properties of the system [34].
@@
The moiré periodicity for a given twist angle $\theta$ decreases as the strain increases. The above condition does not, however, guarantee that both moiré patterns (with and without strain) would be aligned because their orientation may differ. This can be important in superlattice configurations made up of three (or more) layers with a lattice mismatch, and where only two layers are relatively rotated. In general, the angle $\alpha$ of the moiré pattern orientation, with respect to the non-strain case, reads [34]
@@
A relevant example of the importance of matching the periodicity and orientation condition occurs in heterostructures of TBG/hBN, in which hBN acts as a substrate of TBG [126, 124, 31]. This situation can be described by considering three equal graphene monolayers, but with a biaxial strain of magnitude $\epsilon_{b}\sim 1-a_{g}/a_{h}=0.016$ in the bottom layer [24], in order to account for the lattice mismatch between graphene ($a_{g}=2.46\;\textrm{Å}$) and hBN ($a_{h}=2.50\;\textrm{Å}$). Since the moiré pattern remains hexagonal with twist and biaxial strain, the orientation match between the graphene/hBN moiré and the TBG moiré occurs when the angle $\alpha$ in Eq. (64) is $60^{\circ}$ (or integers values of it), so that $\cos\alpha=\pm 1/2$. Combining this with the equal periodicity condition of Eq. (63), one finds the moiré alignment conditions [76, 77, 24, 34]
@@ III.1.4 Shear and biaxial strain
The possible geometries of the strained moiré pattern, as given by the transformation (68), are analogous to those that result from uniaxial heterostrain. One can actually state this analogy precisely by noting that both cases are equivalent by the correspondence [34]
@@
This allows one to readily obtain the geometrical properties of the strained moiré patterns due to shear and biaxial strain from those previously analyzed for uniaxial heterostrain (Section III.1.1). In particular, one can directly replace the above correspondence in Eqs. (57) and (58) and obtain the needed combination of shear and biaxial strain in order to obtain equal-length moiré vectors with an angle $\beta$ between them. It should be noted that this analogy between different kinds of strain is exact, it does not just alludes to an equivalence of periodicity or orientation. Indeed, the above correspondence leads to equivalent uniaxial and shear+biaxial strain tensors, so that each lattice is exactly strained in the same way. Thus, one can think of uniaxial heterostrain as a particular combination of biaxial and shear strain [60, 34]. As the Poisson’s ratio is generally fixed by the properties of the material, this analogy necessarily fixes the ratio between shear and biaxial strain magnitude
@@
Figure 4: Variation of the real space moiré vectors lengths $A^{\prime}_{1}=\left|\mathbf{G}_{1}^{R}\right|$ and $A^{\prime}_{2}=\left|\mathbf{G}_{2}^{R}\right|$, as a function of the twist angle and the strain magnitude, for the cases of two homobilayer ($\mathrm{WSe_{2}-WSe_{2}}$) and heterobilayer ($\mathrm{WSe_{2}}-\mathrm{MoSe_{2}}$) hexagonal lattices. The configuration is such that the top layers is only twisted and the bottom layer is only strained. Panel (a) shows the moiré lengths $A^{\prime}_{1},A^{\prime}_{2}$ (equal) under biaxial strain magnitude $\epsilon_{b}$; dotted ellipses indicate the regions of moiré lenghts $A^{\prime}_{1},A^{\prime}_{2}=10\,\mathrm{nm}.$ Panel (b) shows the moiré lengths $A^{\prime}_{1}$ (top) and $A^{\prime}_{2}$ (bottom) under uniaxial strain magnitude $\epsilon_{u}$ at fixed direction $\phi=0^{\circ}$. The spots labeled as I, II and III indicate the unidimensional, square and hexagonal moiré geometries, respectively (cf. Section III.4). The points $\nu_{1}$ and $\nu_{2}$ in the heterobilayer case indicate the vertexes of the divergence curves (unidimensional patterns). Adapted under the terms of the $C$ BY license from Ref. [60]. Copyright (2023) The author(s).
@@ III.2 Hexagonal heterobilayers
The analysis of the previous section can be directly generalized to the case of two hexagonal layers with different lattice constant. We refer to these as heterobilayers. Since then the non-deformed lattice vectors in each lattice are different, one cannot determine the moiré vectors through a transformation $\mathbf{T}$ as in Eq. (46). However, one can still define a generalized transformation $\mathbf{T}_{h}$ by accounting the lattice mismatch with a biaxial strain $\mathcal{E}_{b}=\mathbb{I}\epsilon_{b}$ [27, 60]. Since the moiré pattern only depends on relative deformations, it is convenient to introduce a biaxial strain tensor $\pm\mathcal{E}_{b}/2$ so that the moiré vectors in heterobilayers can be determined as $\mathbf{G}_{i}=\mathbf{T}_{h}\mathbf{b}_{i}$, where
@@
Figure 4 shows the variation in the length of the real space moiré vectors $A^{\prime}_{1},A^{\prime}_{2}=\left|\mathbf{G}_{1}^{R}\right|,\left|\mathbf{G}_{2}^{R}\right|$, as a function of the twist angle and the strain magnitude (biaxial or uniaxial), for the cases of hexagonal homobilayers and heterobilayers. The results correspond to $\mathrm{WSe_{2}}$ and $\mathrm{MoSe_{2}-WSe_{2}}$ homobilayers and heterobilayers, with lattice constant $a_{\mathrm{MoSe_{2}}}=0.329\,\mathrm{nm}$ and $0.4\%$ smaller for $a_{\mathrm{WSe_{2}}}$, and Poisson’s ratios $\nu_{\mathrm{MoSe_{2}}}=0.23$, $\nu_{\mathrm{WSe_{2}}}=0.19$. The results correspond to a particular non-symmetric configuration in which the upper layer is only twisted while the lower layer is only strained.
@@
Figure 5: Variation of the real-space angle $\beta_{R}$ between the moiré vectors (top) and the moiré area $M=\left|\mathbf{G}_{1}^{R}\times\mathbf{G}_{2}^{R}\right|$ (bottom), as a function of the twist angle and the strain magnitude, for the cases of two homobilayer ($\mathrm{WSe_{2}-WSe_{2}}$) and heterobilayer ($\mathrm{WSe_{2}}-\mathrm{MoSe_{2}}$) hexagonal lattices (same configuration as in Figure 4). Panel (a) shows the case of uniaxial strain with magnitude $\epsilon_{u}$ at fixed direction $\phi=0^{\circ}$. Panel (b) shows the case of shear strain with magnitude $\epsilon_{s}$ at fixed direction $\phi=0^{\circ}$. In both cases, the spots labeled as I, II and III indicate the unidimensional, square and hexagonal moiré geometries (see also Section III.4). Adapted under the terms of the $C$ BY license from Ref. [60]. Copyright (2023) The author(s).